OpenStudy (anonymous):

Given the function T(z) = z – 6, find T(–4). A. 2 B. 10 C. –10 D. –2

OpenStudy (unklerhaukus):

T( z) = ( z) – 6 T(–4) = (-4) - 6 = -4 - 6

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

what did you get ?